# The Honeycomb Conjecture in normed planes and an alpha-convex variant of a theorem of Dowker

### v2

@@ Abstract.
Partially supported by the ERC Advanced Grant “ERMiD”, the National Research, Development and Innovation Office, NKFI, K-147544 grant, and BME Water Sciences & Disaster Prevention TKP2020 Institution Excellence Subprogram, grant no. TKP2020 BME-IKA-VIZ
@@ 1. Introduction
Clearly, it is a natural question to ask if a similar statement holds for tilings in any normed plane. The aim of this paper is to investigate this problem. To do it, we recall a few well-known facts about normed planes. In particular, we recall that every origin-symmetric convex disk $M$ (i.e. every compact, convex set $M$ with nonempty interior and satisfying $M=-M$) is the unit disk of a normed plane, and the unit disk of a normed plane is an origin-symmetric convex disk. In the paper, for a normed plane $\mathcal{M}$, we denote the unit disk of $\mathcal{M}$ by $M$, and the norm of a point $p\in\mathcal{M}$ by $||p||_{M}$. For convenience, we imagine that $\mathcal{M}$ also has an underlying Euclidean structure, and denote the Euclidean norm of $p$ by $||p||$.
@@
Is it true that in any normal, convex tiling in any normed pane $\mathcal{M}$, with unit area cells, the average perimeter of a cell, if it exists, is minimal for a hexagonal tiling?
@@
Let $\mathcal{T}$ be a convex, normal tiling in the normed plane $\mathcal{M}$. For any $R>0$, let $\mathcal{T}(R)$ denote the family of cells of $\mathcal{T}$ contained in $R\mathbf{B}^{2}$. Let $\alpha>0$. We define the lower average $\alpha$th powered perimeter of a cell of $\mathcal{T}$ as the quantity
@@
Similarly, we say that it satisfies the log-honeycomb (or 00-honeycomb) property if the same holds for the lower average log-perimeter of a cell of $\mathcal{T}$.
@@
Let $\alpha\in(0,\infty)$. We say that a convex disk $K$ satisfies the $\alpha$-Dowker property if the sequence $\{A_{K}^{\alpha}(n)\}$ is convex. Furthermore, we say that it satisfies the log-Dowker (or 00-Dowker) property if the sequence $\{\log A_{K}(n)\}$ is convex.
@@
holds for any $3\leq m<6<n$. Similarly, we say that $K$ satisfies the weak log-Dowker (or weak 00-Dowker) property if
@@
First, we observe that Theorem 1.9, combined with Theorem 1.5 readily implies Theorem 1.4. Thus, we need only prove Theorem 1.9. Throughout this section, we deal with a given convex, normal tiling $\mathcal{T}$ in a normed plane $\mathcal{M}$. We permit the angles of cells to be straight angles; due to this, we can regard $\mathcal{T}$ as an edge-to-edge tiling, in which every edge of every cell belongs to exactly one more cell. During our investigation, $\mathcal{T}(R)$ denotes the family of cells of $\mathcal{T}$ contained in $R\mathbf{B}^{2}$, and for every $C\in\mathcal{T}$, $v(C)$ denotes the number of sides of $C$. We start with collecting some elementary observations and known facts about tilings and normed planes.
@@
[…]
@@
Now we are ready to prove Theorem 1.9. For simplicity, we assume that $M_{iso}$ satisfies the $\alpha$-Dowker property for some $\alpha\in(0,\infty)$; for the case $\alpha=0$ a straightforward modification of our argument can be applied. From now on we also assume that every cell of $\mathcal{T}$ has unit area. We need to consider the following quantity.
@@
where $A_{M_{iso}}(v(C))$ denotes the area of a minimum area polygon with $v(C)$ sides, circumscribed about $M_{iso}$ (see Theorem 1.5). By our assumptions, the sequence $\{\left(A_{M_{iso}}(n)\right)^{\alpha}\}$ is convex, and by Lemma 2.4, $\overline{v}(\mathcal{T})\leq 6$. Thus,
@@ 3.1. An algorithm to check the $\alpha$-honeycomb property of a polygonal norm
In this subsection we give a simple algorithm that checks if the isoperimetrix of a polygonal norm satisfies the $\alpha$-Dowker, or weak $\alpha$-Dowker property for some given $\alpha\in[0,\infty)$. We assume that $\alpha>0$, and note that our argument can be modified in a natural way for the log-Dowker or weak log-Dowker property.
@@ 3.2. Norms where the unit disk is a regular polygon
During the proof we show (see Lemma 3.7) that if $k\geq 4$, then $A_{2k-2}(P_{k})$ is a convex combination of $A_{2k-1}(P_{k})$ and $A_{2k-3}(P_{k})$, implying that in this case $P_{k}$ does not satisfy the $\alpha$-Dowker property for any $\alpha<1$.
@@
Clearly, every side of $Q$ intersects $P$. Suppose that there is a side $S$ of $Q$ such that the midpoint of $S$ does not belong to $P$. Let $q$ be the point of $P\cap S$ closest to the midpoint of $S$. Then slightly rotating the sideline of $S$ around $q$ in a suitable direction yields a circumscribed polygon whose area is strictly less than $A_{n}(P)$, a contradiction.
@@
Let $Q$ be a minimum area $m$-gon circumscribed about $P_{k}$ with the smallest number of sides that contain only a vertex of $P_{k}$. By Lemma 3.5 and Remark 3.6, we can assume that if $m$ is even, then every side of $Q$ contains a side of $P$, and otherwise at most one side of $Q$ contains only a vertex of $P$. Let the sides of $P_{k}$ be denoted as $S_{1},S_{2},\ldots,S_{2k}$ in counterclockwise order in $\operatorname{bd}(P_{k})$.
@@
Subcase 1.2, there is a unique side of $Q$ that contains only a vertex of $P_{k}$. Observe that by our assumptions, in this case $n\geq 5$ is odd. From this, by Lemma 3.5, it follows that $k\geq n\geq 5$. Let the sides of $Q$ be denoted by $E_{1},E_{2},\ldots,E_{n}$ in counterclockwise order. For $i=2,\ldots,n$, let $2\alpha_{i}$ denote the turning angle of $Q$ at the common vertex of $E_{i-1}$ and $E_{i}$, and let $2\alpha_{1}$ denote the turning angle at the common vertex of $E_{n}$ and $E_{1}$. Without loss of generality, let $E_{1}$ be the side of $Q$ that contains only a vertex of $P_{k}$. Then, for any $i$ with $3\leq i\leq n$, $\alpha_{i}$ is an integer multiple of $\frac{\pi}{2k}$, and $\sum_{i=1}^{n}\alpha_{i}=\pi$. Since $P_{k}$ is circumscribed about $\mathbf{B}^{2}$, the distance of the line of $E_{1}$ from $o$ is strictly greater than $1$. Let us translate the line of $E_{1}$ towards $o$ until its distance is $1$, and let $Q^{\prime}$ denote the convex $n$-gon obtained in this way (see Figure 2). Then
@@
Assume that $\alpha_{2}-\alpha_{1}<\frac{\pi}{2k}$. Then $\frac{\pi}{4}<\alpha_{1}\leq\frac{\pi}{4}+\frac{\pi}{4k}\leq\alpha_{2}<\frac{\pi}{4}+\frac{\pi}{2k}$. Since $\sum_{i=_{3}}^{n}\tan\alpha_{i}=\frac{\pi}{2}-\frac{\pi}{2k}$ and $5\leq n\leq k$, for at least one value of $i$ we have
$\alpha_{i}\geq\frac{1}{n-2}\left(\frac{\pi}{2}-\frac{\pi}{2k}\right)=\frac{(k-1)\pi}{2k(n-2)}>\frac{\pi}{2k}.$
Without loss of generality, let $\alpha_{3}>\frac{\pi}{2k}$. Since $\alpha_{3}$ is an integer multiple of $\frac{\pi}{2k}$, this yields that $\alpha_{3}\geq\frac{\pi}{k}$. Set $\alpha_{2}^{\prime}=\alpha_{2}+\frac{\pi}{2k}$, $\alpha_{3}^{\prime}=\alpha_{3}-\frac{\pi}{2k}$. Then, by the convexity of the $x\mapsto\tan x$ function, we have that $\sum_{i=1}^{n}\tan\alpha_{i}>\alpha_{1}+\alpha_{2}^{\prime}+\alpha_{3}^{\prime}+\sum_{i=4}^{n}\tan\alpha_{i}$. Here $\alpha_{2}^{\prime}-\alpha_{1}>\frac{\pi}{2k}$, and applying the argument in the previous case we obtain a convex $n$-gon $Q^{\prime\prime}$ circumscribed about $P_{k}$ such that $\operatorname{area}(Q^{\prime\prime})<\operatorname{area}(Q^{\prime})$, a contradiction.
@@
Subcase 2.1, every side of $Q$ contains only a vertex of $P_{k}$. Let these vertices of $P_{k}$ be $p_{i},p_{j},p_{l}$, where we can assume that $1\leq i<j<l\leq 2k$. Since these vertices are the midpoints of the corresponding sides of $Q$, $\operatorname{conv}\{p_{i},p_{j},p_{l}\}$ is the midtriangle of $Q$ and hence, the sides of this triangle are parallel to the sides of $Q$. Since $Q$ is circumscribed about $P_{k}$, it follows that $j-i,l-j$ and $i+2k-l$ differ by at most one. Thus, by an elementary computation we obtain that if $Q$ is a minimal area triangle circumscribed about $P_{k}$, and $\operatorname{bd}(Q)$ contains no side of $T$, then, with the notation $2k=3q+r$, where $q,r$ are nonnegative integers and $r\leq 2$,
@@
Subcase 2.2, exactly one side of $Q$ contains a side of $P_{k}$. Let the sides of $Q$ be $E_{1},E_{2},E_{3}$, and assume that $[p_{2k},p_{1}]\subset E_{1}$, and $p_{i}\in E_{2}$ and $p_{j}\in E_{3}$. Since the latter two vertices are midpoints of $E_{2}$ and $E_{3}$, respectively, we have that $[p_{i},p_{j}]$ is parallel to $[p_{2k},p_{1}]$, implying that $j=2k+1-i$. Let $q$ denote the midpoint of $E_{1}$, and recall that $q\in[p_{2k},p_{1}]$. Note that $\operatorname{conv}\{q,p_{i},p_{j}\}$ is the midtriangle of $Q$, and thus, the area of $Q$ does not depend on the position of $q$. This yields that if $\operatorname{conv}\{q,p_{i},p_{j}\}$ is the midtriangle of a triangle circumscribed about $P_{k}$, then so is $\operatorname{conv}\{p_{1},p_{i},p_{j}\}$ or $\operatorname{conv}\{p_{2k},p_{i},p_{j}\}$. Thus, we may apply the argument from Subcase 2.1 and obtain the same quantities for $\operatorname{area}(Q)$ as in (3.4).
@@
for any $q\geq 3$ and $60\leq r<$ 6. Thus, (3.5) holds for any $k\geq 9$. We also observe that since $A(k,n)=\operatorname{area}(P_{k})$ for any $n\geq 2k$, (3.5) clearly holds if $k=2$ or $k=3$. Thus, we are left with the cases $k=4,5,6,7,8$, for which a direct computation of the two sides of (3.5) for any $3\leq m<6<n\leq 2k$, using a Maple 18.00 software, yields the assertion. ∎

### v4

@@ Abstract.
Partially supported by the ERC Advanced Grant “ERMiD”, the National Research, Development and Innovation Office, NKFI, K-147544 grant, and project TKP2021-NVA-09 provided by the Ministry of Innovation and Technology of Hungary.
@@ 1. Introduction
Clearly, it is a natural question to ask if a similar statement holds for tilings in any normed plane. The aim of this paper is to investigate this problem. To do it, we recall a few well-known facts about normed planes. In particular, we recall that every origin-symmetric convex disk $M$ (i.e. every compact, convex set $M$ with nonempty interior and satisfying M=MM=-M) is the unit disk of a normed plane, and the unit disk of a normed plane is an origin-symmetric convex disk. In the paper, for a normed plane $\mathcal{M}$, we denote the unit disk of $\mathcal{M}$ by $M$, and the norm of a point $p\in\mathcal{M}$ by $p||_{M}$. For convenience, we imagine that $\mathcal{M}$ also has an underlying Euclidean structure, and denote the Euclidean norm of $p$ by $p||$.
@@
Is it true that in any normal, convex tiling in any normed plane $\mathcal{M}$, with unit area cells, the average perimeter of a cell, if it exists, is minimal for a hexagonal tiling?
@@
[…]
@@
For any normed plane $\mathcal{M}$ there is a hexagonal tiling $\mathcal{T}_{hex}$ of $\mathcal{M}$ with unit area tiles such that for any convex, normal tiling $\mathcal{T}$ of $\mathcal{M}$ with unit area tiles, we have
@@
Similarly, we say that it satisfies the log-honeycomb (or 0-honeycomb) property if the same holds for the lower average log-perimeter of a cell of $\mathcal{T}$.
@@
Let $\alpha\in(0,\infty)$. We say that a convex disk $K$ satisfies the $\alpha$-Dowker property if the sequence $\{A_{K}^{\alpha}(n)\}$ is convex. Furthermore, we say that it satisfies the log-Dowker (or 0-Dowker) property if the sequence $\{\log A_{K}(n)\}$ is convex.
@@
holds for any $3\leq m<6<n$. Similarly, we say that $K$ satisfies the weak log-Dowker (or weak 0-Dowker) property if
@@
First, we observe that Theorem 1.9, combined with Theorem 1.5 readily implies Theorem 1.4. Thus, we only need to prove Theorem 1.9. Throughout this section, we deal with a given convex, normal tiling $\mathcal{T}$ in a normed plane $\mathcal{M}$. We permit the angles of cells to be straight angles; due to this, we can regard $\mathcal{T}$ as an edge-to-edge tiling, in which every edge of every cell belongs to exactly one more cell. During our investigation, $\mathcal{T}(R)$ denotes the family of cells of $\mathcal{T}$ contained in $R\mathbf{B}^{2}$, and for every $C\in\mathcal{T}$, $v(C)$ denotes the number of sides of $C$. We start with collecting some elementary observations and known facts about tilings and normed planes.
@@
We define the upper average number of sides $\overline{v}(\mathcal{T})$ of a cell of $\mathcal{T}$ in the same way, replacing $\liminf$ by $\limsup$. If $\underline{v}(\mathcal{T})=\overline{v}(\mathcal{T})$, we call this quantity the average number of sides of a cell of $\mathcal{T}$.
@@
Now we are ready to prove Theorem 1.9. We assume that $M_{iso}$ satisfies the weak $\alpha$-Dowker property for some $\alpha\in(0,\infty)$; for the case $\alpha=0$ a straightforward modification of our argument can be applied. From now on we also assume that every cell of $\mathcal{T}$ has unit area. We need to consider the following quantity.
@@
where $A_{M_{iso}}(v(C))$ denotes the area of a minimum area polygon with $v(C)$ sides, circumscribed about $M_{iso}$ (see Theorem 1.5). By Lemma 2.4, $\overline{v}(\mathcal{T})\leq 6$. Now, we leave to the reader that the fact that $A_{M_{iso}}(v(C))$ is weak $\alpha$-Dowker and monotone decreasing yields the following: If $v_{1},\ldots,v_{k}\geq 3$ are real numbers and $0\leq\tau_{1},\tau_{2}\ldots,\tau_{k}\leq 1$ are weights satisfying $\sum_{i=1}^{k}\tau_{i}=1$ and $\sum_{i=1}^{k}\tau_{k}v_{k}\leq 6$, then $\sum_{i=1}^{k}\tau_{i}\left(A_{M_{iso}}(v_{i})\right)^{\alpha}\geq\left(A_{M_{iso}}(6)\right)^{\alpha}$. Based on this, it follows that
@@ 3.1. An algorithm to check the $\alpha$-honeycomb property of a polygonal norm
There is no straightforward way to check whether in a specific normed plane the honeycomb conjecture is true or not. Nevertheless, the sufficient condition in Theorem 1.9 can be checked using a suitable algorithm. In this subsection we give a simple algorithm that checks if the isoperimetrix of a polygonal norm satisfies the $\alpha$-Dowker, or weak $\alpha$-Dowker property for some given $\alpha\in[0,\infty)$. We assume that $\alpha>0$, and note that our argument can be modified in a natural way for the log-Dowker or weak log-Dowker property.
@@ 3.2. Norms where the unit disk is a regular polygon
Consider a normed plane $\mathcal{M}$ whose unit disk is a regular $(2k)$-gon with $k=4,5$ or $7$. A possible interpretation of Theorem 3.3 is that in this case to minimize the ‘average’ perimeter it is not worth using ‘optimal’ hexagons. Nevertheless, it seems unlikely that one can fit ‘optimal’ $k$-gons with more than one different values of $k$ to tile the plane. This observation might indicate that these planes may also satisfy the honeycomb property.
@@
According to Lemma 3.7, if $k\geq 4$, then $A_{2k-2}(P_{k})$ is a convex combination of $A_{2k-1}(P_{k})$ and $A_{2k-3}(P_{k})$, implying that in this case $P_{k}$ does not satisfy the $\alpha$-Dowker property for any $\alpha<1$.
@@
Clearly, every side of $Q$ intersects $P$. Suppose that there is a side $S$ of $Q$ such that the midpoint of $S$ does not belong to $P$. Let $q$ be the point of $P\cap S$ closest to the midpoint of $S$. Then slightly rotating the sideline of $S$ around $q$ in a suitable direction yields a circumscribed polygon whose area is strictly smaller than $A_{n}(P)$, a contradiction (see Figure 2).
Figure 2. Slightly rotating the side $S$ of $Q$ around $q$. The rotated copy is shown with a dashed line. Note that if the longer part of $S$ moves towards $P$, then the area of $Q$ decreases.
@@
Let $Q$ be a minimum area $n$-gon circumscribed about $P_{k}$ with the smallest number of sides that contain only a vertex of $P_{k}$. By Lemma 3.5 and Remark 3.6, we can assume that if $n$ is even, then every side of $Q$ contains a side of $P_{k}$, and otherwise at most one side of $Q$ contains only a vertex of $P_{k}$. Let the sides of $P_{k}$ be denoted as $S_{1},S_{2},\ldots,S_{2k}$ in counterclockwise order in $\operatorname{bd}(P_{k})$.
@@
Subcase 1.2, there is a unique side of $Q$ that contains only a vertex of $P_{k}$. Observe that by our assumptions, in this case $n\geq 5$ is odd. By Lemma 3.5, it follows that the sum of the angles $\alpha,\beta$ of $Q$ on the above side of $Q$ satisfies $\alpha+\beta<\pi$, and thus, at least one half of the sides of $P_{k}$ do not lie in $\operatorname{bd}(Q)$, implying that $k\geq n\geq 5$. Let the sides of $Q$ be denoted by $E_{1},E_{2},\ldots,E_{n}$ in counterclockwise order. For $i=2,\ldots,n$, let $2\alpha_{i}$ denote the turning angle of $Q$ at the common vertex of $E_{i-1}$ and $E_{i}$, and let $2\alpha_{1}$ denote the turning angle at the common vertex of $E_{n}$ and $E_{1}$. Without loss of generality, let $E_{1}$ be the side of $Q$ that contains only a vertex of $P_{k}$. Then, for any $i$ with $3\leq i\leq n$, $\alpha_{i}$ is an integer multiple of $\frac{\pi}{2k}$, and $\sum_{i=1}^{n}\alpha_{i}=\pi$. Since $P_{k}$ is circumscribed about $\mathbf{B}^{2}$, the distance of the line of $E_{1}$ from $o$ is strictly greater than $1$. Let us translate the line of $E_{1}$ towards $o$ until its distance is $1$, and let $Q^{\prime}$ denote the convex $n$-gon obtained in this way (see Figure 3). Then
@@
Assume that $\alpha_{2}-\alpha_{1}<\frac{\pi}{2k}$. Then $\frac{\pi}{4}<\alpha_{1}\leq\frac{\pi}{4}+\frac{\pi}{4k}\leq\alpha_{2}<\frac{\pi}{4}+\frac{\pi}{2k}$. Since $\sum_{i=3}^{n}\alpha_{i}=\frac{\pi}{2}-\frac{\pi}{2k}$ and $5\leq n\leq k$, for at least one value of $i$, with $3\leq i\leq n$, we have
$\alpha_{i}\leq\frac{1}{n-2}\left(\frac{\pi}{2}-\frac{\pi}{2k}\right)\leq\frac{\pi}{6}-\frac{\pi}{6k}.$
Without loss of generality, let $\alpha_{3}$ satisfy this inequality. Then, as $k\geq 5$, $\alpha_{1}-\alpha_{3}>\frac{\pi}{4}-\left(\frac{\pi}{6}-\frac{\pi}{6k}\right)>\frac{\pi}{2k}$. Thus, setting $\alpha_{1}^{\prime}=\alpha_{1}-\frac{\pi}{2k}$, $\alpha_{3}^{\prime}=\alpha_{3}+\frac{\pi}{2k}$, by the convexity of the $x\mapsto\tan x$ function, we have that $\sum_{i=1}^{n}\tan\alpha_{i}>\tan\alpha_{1}^{\prime}+\tan\alpha_{2}+\tan\alpha_{3}^{\prime}+\sum_{i=4}^{n}\tan\alpha_{i}$. Then, applying the argument in the previous case we obtain a convex $n$-gon $Q^{\prime\prime}$ circumscribed about $P_{k}$ such that $\operatorname{area}(Q^{\prime\prime})<\operatorname{area}(Q^{\prime})$, a contradiction.
@@
Subcase 2.1, every side of $Q$ contains only a vertex of $P_{k}$. Let these vertices of $P_{k}$ be $p_{i},p_{j},p_{l}$, where we can assume that $1\leq i<j<l\leq 2k$. Since these vertices are the midpoints of the corresponding sides of $Q$, $\operatorname{conv}\{p_{i},p_{j},p_{l}\}$ is the midtriangle of $Q$ and hence, the sides of this triangle are parallel to the sides of $Q$. Since $Q$ is circumscribed about $P_{k}$, its sidelines support $P_{k}$ at the corresponding vertices of $P_{k}$. Characterizing which diagonals of $P_{k}$ are parallel to some supporting lines at $p_{i}$, $p_{j}$ and $p_{k}$, respectively, we obtain that ji,ljj-i,l-j and i+2kli+2k-l differ by at most one. Thus, by an elementary computation we have that if $Q$ is a minimal area triangle circumscribed about $P_{k}$, and $\operatorname{bd}(Q)$ contains no side of $T$, then, with the notation $2k=3q+r$, where $q,r$ are nonnegative integers and $r\leq 2$,
@@
Subcase 2.2, exactly one side of $Q$ contains a side of $P_{k}$. Let the sides of $Q$ be $E_{1},E_{2},E_{3}$, and assume that $p_{2k},p_{1}]\subset E_{1}$, and $p_{i}\in E_{2}$ and $p_{j}\in E_{3}$. Since the latter two vertices are midpoints of $E_{2}$ and $E_{3}$, respectively, we have that $p_{i},p_{j}]$ is parallel to $p_{2k},p_{1}]$, implying that j=2k+1ij=2k+1-i. Let $q$ denote the midpoint of $E_{1}$, and recall that $q\in[p_{2k},p_{1}]$. Note that $\operatorname{conv}\{q,p_{i},p_{j}\}$ is the midtriangle of $Q$, and thus, the area of $Q$ does not depend on the position of $q$. This yields that as $\operatorname{conv}\{q,p_{i},p_{j}\}$ is the midtriangle of a triangle circumscribed about $P_{k}$, then, moving $q$ along the side $p_{1},p_{2k}$, we obtain that $\operatorname{conv}\{p_{1},p_{i},p_{j}\}$ or $\operatorname{conv}\{p_{2k},p_{i},p_{j}\}$ is also the midtriangle of a triangle circumscribed about $P_{k}$, and the area of this triangle is equal to $\operatorname{area}(Q)$. Thus, we may apply the argument from Subcase 2.1 and obtain the same quantities for $\operatorname{area}(Q)$ as in (3.4).
@@
for any $q\geq 3$ and $0\leq r<6$. Thus, (3.5) holds for any $k\geq 9$. We also observe that since $A(k,n)=\operatorname{area}(P_{k})$ for any $n\geq 2k$, (3.5) clearly holds if $k=2$ or $k=3$. Thus, we are left with the cases $k=4,5,6,7,8$, for which a direct computation of the two sides of (3.5) for any $3\leq m<6<n\leq 2k$, using a Maple 18.00 software, yields the assertion. ∎